OpenStudy (anonymous):

In 2000, the circulation of a local newspaper was 3,250. In 2001, its circulation was 3,640. In 2002, the circulation was 4,100. a. Find the percent of increase in the newspaper’s circulation from 2000 to 2001 and from 2001 to 2002. b. Which period had a higher percent of increase, 2000 to 2001, or 2001 to 2002? For questions 18, remember to show all of the steps that you use to solve the problem.

OpenStudy (anonymous):

help please @weedman218 @qwerties @Agent_Sniffles @AccessDenied @A.Avinash_Goutham @AlexPR787 @alainabbyboo22 @Alyx_da_Boss @TuringTest @TheViper @TheWriter @technopanda13 @toristar9504 @teddygirl @NoelGreco @brinethery @NikhilRam @natasha.aries @Newnew @niksva @natashaduncanw @BrennaM @babystar16 @bmelyk @batmanismyhero12 @zepdrix @campbell_st @velociraptor @ChmE @Cemantha @Chris2332 @courtneycara17 @Chipper10 @dmezzullo @Daish96 @Dwade03 @Dawayne @damien @dopek.i.d.s @myininaya @mahmit2012 @marcoduuuh @Miss_Deva @msumner @momoluvgd @qwerties @Hero @him1618 @hunterhoff @hitten101 @Hollis @eliassaab @EsmeraldaT @Exvikki @emorainbow @e.g.1 @JakeV8 @jasonxx @JazzyPoowh @JerJason @JetWave600 @kymber @kelly226 @KJ98 @kaiteDID @Kris5 @2yoututu11 @robtobey @radar @rebecca1233 @RioAnne @richyw @Reppinthe80 @iheartfood @ivanmlerner @InsanelyChaotic @imuruki @IamElaine8 @Ice_Cream @phi @Preetha @pkjha3105 @Plutog55 @panchtatvam @princemartin @LoveYou*69 @lolly @LilVeroDez13 @LonelyandForgotten @LukeBlueFive @lollypopabby43 @Xishem @Xero1999 @x-Jamie-x @velociraptor @gerryliyana @Grim0519 @gljay

OpenStudy (anonymous):

@SUpAmArIo @swissgirl @Susan9419 @SUpAmArIo @Swageraid32

OpenStudy (anonymous):

@oldrin.bataku @Omniscience @ovoxodrake10 @OliveBandit00

OpenStudy (anonymous):

@Fools101 @Frostbite @FirstFrostByte @Figureskater120

OpenStudy (anonymous):

i appreciate you taggin me to help you but maybe i can help later because i am working on this major project that is 65% of my grade so as you can see..im sort of under a lot of stress right now..

OpenStudy (anonymous):

((3640-3250)/3250 ) * 100 n same for the other part

OpenStudy (anonymous):

compare percentages n see which is greater

OpenStudy (kymber):

It's generally not a good idea to tag this many people in your question. Just tag one or two people that you know can answer it.

OpenStudy (lukebluefive):

a. Increase from 2000 to 2001: \[\frac{3640}{3250} \times 100\% = ?\%\] Increase from 2001 to 2002: \[\frac{4100}{3640} \times 100\% = ?\%\] b. Compare answers. c. Show all work.

OpenStudy (anonymous):

You don't need to call the entire website here to solve your question! People look at the questions and answer when they think they can help, if you have a preference for someones way of explaining call this person but not everybody.

OpenStudy (anonymous):

a) From 2000 to 2001, percent increase = 3640 − 3250 3250 = 0.12 = 12 100 = 12% From 2001 to 2002, percent increase = 4100 − 3640 3640 =Ö 0.13 = 13 100 = 13% b) The period 2001 to 2002 had the higher percent increase

OpenStudy (anonymous):

@AwesomePowers=)!